A solution of water and alcohol is 25% alcohol. If the amount of water is doubled what percent is now alcohol?
Answer by Theo(13342)

the way i see this problem is as follows:

you have a solution of water and alcohol.
the solution is 3 parts water and 1 part alcohol.
the percent of alcohol is 1/4 = .25 = 25%.

if you double the water, they you are adding another 3 parts of water to get:

6 parts water and 1 part alcohol.
the percent of alcohol is 1/7 = .1428571429 = 14.28571429%.

as a test, i assumed 100 liters of solution that is 75% water and 25% alcohol.
that divides into 3 parts water and 1 part alcohol.
that means 75 liters water and 25 liters alcohol.
i doubled the amount of water to get 150 liters water and 25 liters alcohol in a total solution of 175 liters.
the percent alcohol became 25/175 = .1428571429 = 14.28571429%.

method looks good.
solution is 14.29% alcohol rounded to 2 decimal places.
round your number as required.
the detailed solution is 14.28571429% alcohol.